### v1

@@ 0.3.3 Geometric Analysis of Embedding Spaces
We start with an observation that, given an NLP dataset $\mathcal{Y}$ that contains (a finite number of sentences) $s_{1},\ldots,s_{q}$, and an embedding function $\mathbb{E}:\mathcal{Y}\rightarrow^{m}$, we can define an embedding matrix $\mathcal{X}\in^{q\times m}$, where each row $j$ is given by $\mathbb{E}(s_{j})$. Treating embedded sentences as matrices, rather than as points in the real vector space, makes many computations easier. This motivates the following definition of a hyper-rectangle for $\mathcal{X}$.
Definition 1 (Hyper-rectangle for an Embedding Matrix).
@@
$\mathbb{H}_{\mathcal{X}}\coloneqq\{x\in^{m}:\forall j\in[1,...,m].\>x^{j}\in[\min\nolimits_{i}(\mathcal{X}^{ij}),\max\nolimits_{i}(\mathcal{X}^{ij})]\}.$
where $\min\nolimits_{i}(\mathcal{X}^{ij})$ stands for the minimum value of the $j$-th element $x^{j}$ across all rows $i\in\{0,\ldots,q\}$ in $\mathcal{X}$; similarly for $\max\nolimits_{i}(\mathcal{X}^{ij})$.
Definition 2 (Subspace of the Embedding Space).
Given a NLP dataset $\mathcal{Y}$, an embedding function $\mathbb{E}:\mathcal{Y}\rightarrow^{m}$, and $\mathcal{Y}^{\prime}\subseteq\mathcal{Y}$ (with $\mathcal{Y}^{\prime}=\{s_{1},\ldots,s_{q}\}$), we say a hyper-rectangle $\mathbb{H}_{\mathcal{X}^{\prime}}$ is a subspace of the embedding space ${}^{m}$ around $\mathcal{Y}^{\prime}$ if the matrix $\mathcal{X}^{\prime}\in^{q\times m}$ satisfies the following condition: each $j$th row is given by $\mathbb{E}(s_{j})$.
@@
A computationally efficient way of making hype-rectangles more precise is to rotate them to align to the position of the given points in the embedding space. This motivates us to introduce the Eigenspace rotation.
Eigenspace Rotation.
To construct the tightest possible hyper-rectangle, we define a specific method of eigenspace rotation. As shown in Figure 1 (C and D), our approach is to calculate a rotation matrix $A$ such that the rotated matrix $\mathcal{X}_{\text{rot}}=\mathcal{X}A$ is better aligned with the axes than $\mathcal{X}$, and therefore $\mathbb{H}_{\mathcal{X}_{\text{rot}}}$ has a smaller volume. By a slight abuse of terminology, we will refer to $\mathbb{H}_{\mathcal{X}_{\text{rot}}}$ as the rotated hyper-rectangle, even though strictly speaking, we are rotating the data, not the hyper-rectangle itself. In order to calculate the rotation matrix $A$, we use singular value decomposition [60]. The singular value decomposition of $\mathcal{X}$ is defined as $\mathcal{X}=U\Sigma V^{*}$, where $U$ is a matrix of left-singular vectors, $\Sigma$ is a matrix of singular values and $V^{*}$ is a matrix of right-singular vectors and $\cdot^{*}$ denotes the conjugate transpose. Intuitively, the right-singular vectors $V^{*}$ describe the directions in which $\mathcal{X}$ exhibits the most variance. The main idea behind the definition of rotation is to align these directions of maximum variance with the standard canonical basis vectors. Formally, using $V^{*}$, we can compute the rotation (or change-of-basis) matrix $A$ that rotates the right-singular vectors onto the canonical standard basis vectors $I$, where $I$ is the identity matrix. To do this, we observe that $V^{*}A=I\rightarrow V^{*}=IA^{-1}\rightarrow V^{-1}=A^{-1}\rightarrow V=A$. We thus obtain $\mathcal{X}_{rot}=\mathcal{X}A$ as desired.
All hyper-rectangles constructed and used in this paper are rotated.
Geometric and Semantic Subspaces.
@@
We will finish this section with defining semantic subspaces formally. Given a data set $\mathcal{Y}$, and some defined algorithm for sentence perturbation $\mathcal{P}$, we can form $\mathcal{Y}^{\prime}=\{s^{*}_{1},\ldots,s^{*}_{q}\}$, such that each $s^{*}_{i}=\mathcal{P}(s_{i})$. Intuitively, we want to construct a subspace for $\mathcal{Y}^{\prime}$ as described in Definition 2. However, in the later sections, we will need to refer to different kinds of perturbations (e.g. character, word or sentence level as explained in Section 0.3.1) or different types of perturbations (e.g. insertion, deletion, replacement, etc.) as illustrated in Tables 3 and 4). This motivates the following definitions. Notation $\mathcal{P}_{k}^{t}(s,a)$ will denote a perturbation of kind $k$ and type $t$, applied in the sentence $s$, in a position $a$. As all algorithms we use choose $a$ randomly, we will omit $a$ and write just $\mathcal{P}_{k}^{t}(s)$. When we say that a sentence $s$ has $b$ semantic perturbations of kind $k$, we refer to a set $\mathcal{A}^{k}_{s}=\{\mathcal{P}_{k}^{t_{1}}(s),\mathcal{P}_{k}^{t_{2}}(s),\ldots,\mathcal{P}_{k}^{t_{b}}(s)\}$, for some given choice of $b$ perturbations of types $t_{1},\ldots,t_{b}$. When the choice is unimportant or clear from the context, we may omit mentioning the chosen types of perturbation.
Definition 3 (Semantic Subspace for a Sentence).
Suppose we are given an NLP dataset $\mathcal{Y}=\{s_{1},\ldots,s_{q}\}$, and an embedding function $\mathbb{E}:\mathcal{Y}\rightarrow^{m}$.
Given a sentence $s_{i}$ and a corresponding set $\mathcal{A}^{k}_{s_{i}}$ of its semantic perturbations (of kind $k$), a semantic subspace for $s_{i}$ is a subspace around $\mathcal{Y}^{\prime}=\mathcal{A}^{k}_{s_{i}}$ constructed according to Definition 2.
It will be useful in the later sections to use notation $\mathbb{H}_{\mathcal{X}_{s_{i}}}$ to refer to the hyper-rectangle used in constructing the semantic subspace for $s_{i}$. We will refer to a set of such hyper-rectangles $\mathbb{H}^{*}_{k}=\bigcup_{i=1}^{q}\mathbb{H}^{*}_{\mathcal{X}_{s_{i}}}$, where $s_{i}\in\{s_{1},\ldots,s_{q}\}$ and $k$ is a chosen perturbation kind.
If $\mathbb{H}^{*}_{k}$ is based on perturbations of only one type, we will sometimes use that type as an index to $\mathbb{H}^{*}_{\_}$.
Example 2 (Construction of Semantic Subspaces).
To illustrate this construction, let us consider the sentence $s_{i}=$“Can u tell me if you are a chatbot?”. This sentence is one of $3400$ original sentences of the positive class in the dataset. From this single sentence, we can create six sentences (word-level perturbations), see Table 4. Once these seven sentences are embedded into the vector space, they form the hyper-rectangle $\mathbb{H}_{\mathcal{X}_{s_{i}}}$. By repeating this construction for the remaining $3399$ sentences, we obtain the set of hyper-rectangles $\mathbb{H}^{*}_{word}$ for the dataset.
0.3.4 Training
As outlined in Section 0.2.2, robust training is essential for bolstering the robustness of DNNs; without it, their verifiability would be significantly diminished. This study employs two robust training methods, namely data augmentation and a custom PGD adversarial training, with the goal of discerning the factors contributing to the success of robust training and compare the effectiveness of these methods.
Data Augmentation.
In this training method, we statically generate semantic perturbations at the character, word, and sentence levels before training, which are then added to the dataset. The network is subsequently trained on this augmented dataset using the standard stochastic gradient descent algorithm.
Adversarial Training.
In this training method, we modify the traditional Projected Gradient Descent (PGD) algorithm [85], defined as follows. Given a loss function $\mathcal{L}$, a step size $\gamma\in$ and a starting point $\hat{x}_{0}$ then the output of the PGD algorithm $x(l)$ after $l$ iterations is defined as:
$x(0)$ $=\hat{x}_{0}$
$x(t+1)$ $=\text{proj}_{\mathbb{H}}\Big{[}x(t)+\gamma\cdot\text{sign}(\nabla_{x(t)}\mathcal{L}(x(t),y))\Big{]}$

### v3

@@ 3.4 Working with Embedding Spaces: Our Approach
We start with an observation that, given an NLP dataset $\mathcal{Y}$ that contains a finite set of sentences $s_{1},\ldots,s_{q}$ belonging to the same class, and an embedding function $E:\mathbb{S}\rightarrow^{m}$, we can define an embedding matrix $\mathcal{X}\in^{q\times m}$, where each row $j$ is given by $E(s_{j})$. We will use the notation $x_{i}$ to refer to the $i$th element of the vector $x$, and $\mathcal{X}^{ij}$ to refer to the element in the $i$th row and $j$th column of $\mathcal{X}$. Treating embedded sentences as matrices, rather than as points in the real vector space, makes many computations easier. We can therefore define a hyper-rectangle for $\mathcal{X}$ as follows.
Definition 5 (Hyper-rectangle for an Embedding Matrix)
@@
$\mathbb{H}(\mathcal{X})\coloneqq\{(\min\nolimits_{i=0}^{q}\mathcal{X}^{ij},~{}\max\nolimits_{i=0}^{q}\mathcal{X}^{ij})\mid j\in[1,...,m]\}$
Therefore given an embedding function $E:\mathbb{S}\rightarrow^{m}$, and a set of sentences $\mathcal{Y}=\{s_{1},\ldots,s_{q}\}$, we can form a subspace $\mathbb{H}(E(\mathcal{Y}))$ by constructing the embedding matrix, as described above, and forming the corresponding hyper-rectangle. To simplify the notation, we will omit the application of $E$ and from here on simply write $\mathbb{H}(\mathcal{Y})$.
@@
3.4.1 Exclusion of Unwanted Sentences Via Shrinking
Another concern is that the generated hyper-rectangles may contain sentences from a different class. This would make it unsuitable for verification. In order to exclude all samples from the wrong class, we define a shrinking algorithm $SH(\mathcal{X},\mathcal{Y},c)$ that calculates a new subspace that is a subset of the original hyper-rectangle around $\mathcal{X}$, that only contains embeddings of sentences in $\mathcal{Y}$ that are of class $c$. Of course, to ensure this, the algorithm may have to exclude some sentences of class $c$. The second graph of Figure 3 gives a visual intuition of how this is done.
Figure 3: An example of hyper-rectangle drawn around all points of the same class (a), shrunk hyper-rectangle $\mathbb{H}_{sh}$ that is obtained by excluding all points from the opposite class (b) and clustered hyper-rectangles (c) in 2-dimensions. The red dots represent sentences in the embedding space of one class, while the blue dots are embedded sentences that do not belong to that class.
Formally, for each sentence $s$ in $\mathcal{Y}$ that is not of class $c$, the algorithm performs the following procedure. If $E(s)$ lies in the current hyper-rectangle $(a_{1},b_{1}),\dots,(a_{m},b_{m})$, then for each dimension ${j\in[1,\dots,m]}$ we compute the distance whether $E(s)_{j}$ is closer to $a_{j}$ or $b_{j}$. Without loss of generality, assume $a_{j}$ is closer. We then compute the number of sentences of class $c$ that would be excluded by replacing $a_{j}$ with $E(s)_{j}+\delta$ in the hyper-rectangle where $\delta$ is a small positive number (we use $e^{-100}$). This gives us a penalty for each dimension $j$, and we exclude $s$ by updating the hyper-rectangle in the dimension that minimises this penalty. The idea is to shrink the hyper-rectangle in the dimensions that exclude as few embedded sentences from the desired class $c$ as possible [^6] Note that this algorithm shrinks exactly one dimension by a minimal amount to exclude the unwanted embedded sentence. This choice keeps the algorithm fast while guaranteeing the subspace to retain the highest number of wanted inputs. However, it is not necessarily the best choice for verification: there might be cases where perturbations of the unwanted input are left inside after shrinking and, if the network classifies them correctly, the subspace can never be verified. For large subspaces, our algorithm might render verification unachievable and more clever algorithms should be explored and discussed..
3.4.2 Exclusion of Unwanted Sentences Via Clustering
An alternative approach to excluding unwanted sentences, is to split the dataset up by clustering semantically similar sentences in the embedding space, and then compute the hyper-rectangles around each cluster individually, as shown in the last graph of Figure 3. In this paper we will use the k-means algorithm for clustering. We will use the notation $CL(\mathcal{Y},k)$ to refer to the $k$-clusters formed by applying it to dataset $\mathcal{Y}$. While in our experiments we have found this is often sufficient to exclude unwanted sentences, it is not guaranteed to do so. Therefore, this method is combined with the shrinking algorithm in our experiments.
3.4.3 Eigenspace Rotation
A final alternative and computationally efficient way of reducing the likelihood that the hyper-rectangles will contain embedded sentences of an unwanted class, is to rotate them to better align to the distribution of the embedded sentences of the desired class in the embedding space. This motivates us to introduce the Eigenspace rotation.
To construct the tightest possible hyper-rectangle, we define a specific method of eigenspace rotation. As shown in Figure 1 (C and D), our approach is to calculate a rotation matrix $A$ such that the rotated matrix $\mathcal{X}_{\text{rot}}=\mathcal{X}A$ is better aligned with the axes than $\mathcal{X}$, and therefore $\mathbb{H}(\mathcal{X}_{\text{rot}})$ has a smaller volume. By a slight abuse of terminology, we will refer to $\mathbb{H}(\mathcal{X}_{\text{rot}})$ as the rotated hyper-rectangle, even though strictly speaking, we are rotating the data, not the hyper-rectangle itself. In order to calculate the rotation matrix $A$, we use singular value decomposition [146]. The singular value decomposition of $\mathcal{X}$ is defined as $\mathcal{X}=U\Sigma V^{*}$, where $U$ is a matrix of left-singular vectors, $\Sigma$ is a matrix of singular values and $V^{*}$ is a matrix of right-singular vectors and $\cdot^{*}$ denotes the conjugate transpose. Intuitively, the right-singular vectors $V^{*}$ describe the directions in which $\mathcal{X}$ exhibits the most variance. The main idea behind the definition of rotation is to align these directions of maximum variance with the standard canonical basis vectors. Formally, using $V^{*}$, we can compute the rotation (or change-of-basis) matrix $A$ that rotates the right-singular vectors onto the canonical standard basis vectors $I$, where $I$ is the identity matrix. To do this, we observe that $V^{*}A=I$ implies $V^{*}=IA^{-1}$, which implies $V^{-1}=A^{-1}$, and thus $V=A$. We thus obtain $\mathcal{X}_{rot}=\mathcal{X}A$ as desired. All hyper-rectangles constructed in this paper are rotated.
3.4.4 Geometric and Semantic Subspaces
@@
We will finish this section with defining semantic subspaces formally. We will use $\mathcal{P}_{t}(s)$ to denote an algorithm for generating sentence perturbations of type $t$, applied to an input sentence $s$ in a random position. In the later sections, we will use $t$ to refer to the different types of perturbations illustrated in Tables 4 and 5, e.g. character-level insertion, deletion, replacement. Intuitively, given a single sentence we want to generate a set of semantically similar perturbations and then construct a hyper-rectangle around them, as described in Definition 5.
This motivates the following definitions. Given a sentence $s$, a number $b$, and a type $t$, the set $\mathcal{A}^{b}_{t}(s)=\{\mathcal{P}_{t}(s)\mid i\in[1,b]\}$ is the set of $b$ semantic perturbations of type $t$ generated from $s$. We will use the notation $\mathcal{A}^{b}_{t}(\mathcal{Y})=\bigcup_{s\in\mathcal{Y}}\mathcal{A}^{b}_{t}(s)$ to denote the new dataset generated by creating $b$ semantic perturbations of type $t$ around each sentence.
Definition 6 (Semantic Subspace for a Sentence)
Given an embedding function $E:\mathbb{S}\rightarrow^{m}$, the semantic subspace for a sentence $s$ is the subspace ($\mathbb{H}(\{s\}\cup\mathcal{A}^{b}_{t}(s))$. We will refer to a set of such semantic hyper-rectangles over an entire dataset $\mathcal{Y}$ as $\mathbb{H}_{t}^{b}(\mathcal{Y})=\bigcup_{s\in\mathcal{Y}}\mathbb{H}(\{s\}\cup\mathcal{A}^{b}_{t}(s))$.
Example 2 (Construction of Semantic Subspaces)
To illustrate this construction, let us consider the sentence $s$: “Can u tell me if you are a chatbot?”. This sentence is one of $3400$ original sentences of the positive class in the dataset. From this single sentence, we can create six new sentences using the word-level perturbations from Table 5 to form $\mathcal{A}^{6}_{word}(s)$. Once the seven sentences are embedded into the vector space, they form the hyper-rectangle ($\mathbb{H}(\{s\}\cup\mathcal{A}^{6}_{word}(s))$. By repeating this construction for the remaining $3399$ sentences, we obtain the set of hyper-rectangles $\mathbb{H}_{word}(\mathcal{Y})$ for the dataset.
Given a sentence $s$, we embed each sentence in $\mathcal{A}^{b}_{t}(s)=\{s_{1},\dots,s_{b}\}$ into m obtaining vectors $\mathcal{V}^{b}_{t}(s)=\{v,v_{1},\ldots,v_{b}\}$ where $v_{j}=E(s_{j})$.
3.4.5 Measuring the Quality of Sentence Embeddings
One of our implicit assumptions in the previous sections, is that the embedding function $E$ maps pairs of semantically similar sentences to nearby points in the embedding space. In Section 5.5.2, we will evaluate the accuracy of this assumption using cosine similarity. This metric measures how similar two vectors are in a multi-dimensional space by calculating the cosine of the angle between them:
$\text{CoS}(v_{1},v_{2})=\frac{v_{1}\cdot v_{2}}{\|v_{1}\|\|v_{2}\|}$
3.5 Training
As outlined in Section 2.3, robust training is essential for bolstering the robustness of DNNs; without it, their verifiability would be significantly diminished. This study employs two robust training methods, namely data augmentation and a custom PGD adversarial training, with the goal of discerning the factors contributing to the success of robust training and compare the effectiveness of these methods.
Data Augmentation. In this training method, we statically generate semantic perturbations at the character, word, and sentence levels before training, which are then added to the dataset. The network is subsequently trained on this augmented dataset using the standard stochastic gradient descent algorithm.
Adversarial Training. In this training method, the traditional Projected Gradient Descent (PGD) algorithm [20], is defined as follows. Given a loss function $\mathcal{L}$, a step size $\gamma\in$ and a starting point $\hat{x}_{0}$ then the output of the PGD algorithm $x(l)$ after $l$ iterations is defined as:
$x(0)$ $=\hat{x}_{0}$
$x(t+1)$ $=\text{proj}_{\mathcal{S}}\Big{[}x(t)+\gamma\cdot\text{sign}(\nabla_{x(t)}\mathcal{L}(x(t),y))\Big{]}$
@@ 3.5 Training
In this work, we modify the algorithm to work with custom-defined hyper-rectangles as the subspace. The primary distinction between our customised PGD algorithm and the standard version lies in the definition of the step size. In the conventional algorithm, the step size is represented by a scalar $\gamma\in$ therefore representing a uniform step size in every dimension. In our case the width of $\mathbb{H}$ in each dimension may vary greatly, therefore we transforms $\gamma$ into a vector in m, allowing the step size to vary by dimension. Note that the dot $\cdot$ between $\gamma$ and $sign(\nabla)$ becomes an element-wise multiplication. The resulting customised PGD training seeks to identify the worst perturbations within the custom-defined subspace, and trains the given neural network to classify those perturbations correctly, in order to make the network robust to adversarial inputs in the chosen subspace.
3.6 Choice of Verification Algorithm
As stated earlier, our approach in this study involves the utilization of cutting-edge tools for DNN verification. Initially, we employ ERAN [69], a state-of-the-art abstract interpretation-based method. This choice is made over IBP due to its ability to yield tighter bounds. Subsequently, we conduct comparisons and integrate Marabou [22], a state-of-the-art complete verifier. This enables us to attain the highest verification percentage, maximizing the tightness of the bounds. Additionally, we incorporate $\alpha\beta$-CROWN[42, 43], the best-performing verifier in the VNN-COMP 2024 competition, known for its efficiency in linear bound propagation and branch-and-bound techniques. Moreover, we utilise CORA[55], an abstract interpretation-based verifier that supports zonotope-based verification, allowing us to compare hyper-rectangles and zonotopes in our verification experiments. We will use notation $\mathcal{V}$ to refer to a verifier abstractly.
4 Characterisation of Verifiable Subspaces
  • We start with introducing the metric of generalisability of (verified) subspaces and set-up some baseline experiments.
  • We introduce the problem of the verifiability-generalisability trade-off in the context of geometric subspaces.
  • We show that, compared to geometric subspaces, the use of semantic subspaces helps to find a better balance between generalisability and verifiability.
4.1 Metrics for Understanding the Properties of Embedding Spaces
Let us start with recalling the existing standard metrics used in DNN verification. Recall that we are given an NLP dataset $\mathcal{Y}=\{s_{1},\ldots,s_{q}\}$, moreover we assume that each $s_{i}$ is assigned a correct class from $C=\{c_{1},\ldots,c_{n}\}$. We restrict to the case of binary classification in this paper for simplicity, so we will assume $C=\{c_{1},c_{2}\}$. Furthermore, we are given an embedding function $E:\mathbb{S}\rightarrow^{m}$, and a network $N:^{m}\rightarrow^{n}$. Usually $n$ corresponds to the number of classes, and thus in case of binary classification, we have $N:^{m}\rightarrow^{2}$. An embedded sentence $s\in\mathcal{Y}$ is classified as class $c$ if the value of $c$ in $N(E(s))$ is higher than all other classes.
Accuracy. The most popular metric for measuring the performance of the network is the accuracy of $N$, which is measured as a percentage of sentences in $\mathcal{Y}$ that are assigned to a correct class by $N$. Note that this metric only checks a finite number of points in m given by the dataset.
Verifiability. A verifier $\mathcal{V}$ takes a network $N$, a subspace $\mathcal{S}$ and its designated class $c$ as an input, and outputs 1 if it can prove that $N$ assigns all points in the subspace $\mathcal{S}$ to the class $c$ and 0 otherwise. Consider a verification problem with multiple subspaces $\{\mathcal{S}_{1},\ldots,\mathcal{S}_{l}\}$, where all the points in each subspace should be assigned to a specific class $c_{i}\in\{c_{1},\ldots,c_{n}\}$. In the literature, the most popular metric to measure success rate of the given verifier on $\{\mathcal{S}_{1},\ldots,\mathcal{S}_{l}\}$ is verifiability:
Definition 7 (Verifiability)
Given a set of subspaces $\mathcal{S}_{1},\ldots,\mathcal{S}_{l}$ each assigned to classes $c_{1},\ldots,c_{l}$, then the verifiability is the percentage of such subspaces successfully verified:
$\mathcal{W}(\mathcal{S}_{1},\ldots,\mathcal{S}_{l},c_{1},\ldots,c_{l})=\frac{\sum_{i=0}^{l}\mathcal{V}(N,\mathcal{S}_{i},c_{i})}{l}$
All DNN verification papers that study such problems report this measure. Note that each subspace contains an infinite number of points.
However, suppose we have a subspace $\mathcal{S}$ that verifiably consists only of vectors that are assigned to a class $c$ by $N$. Because of the embedding gap, it is difficult to calculate how many valid unseen sentences outside of $\mathcal{Y}$ will be mapped into $\mathcal{S}$ by $E$, and therefore how much utility there is in verifying $\mathcal{S}$. In an extreme case it is possible to have 100% verifiability and yet the verified subspaces will not contain any unseen sentences.
Generalisability. Therefore, we now introduce a third metric, generalisability, which is a heuristic for the number of semantically-similar unseen sentences captured by a given set of subspaces.
Definition 8 (Generalisability)
Given a set of subspaces $\mathcal{S}_{1},\ldots,\mathcal{S}_{l}$ and a target set of embeddings $V$ the generalisability of the subspaces is measured as the percentage of the embedded vectors that lie in the subspaces:
$\mathcal{G}(V,\mathcal{S}_{1},\ldots,\mathcal{S}_{l})=\frac{|V\cap\bigcup_{i=1}^{l}\mathcal{S}_{i}|}{|V|}$
In this paper we will generate the target set of embeddings $V$ as $\bigcup_{s\in\mathcal{Y}}\mathcal{V}^{b}_{t}(s)$ where $\mathcal{Y}$ is a dataset, $t$ is the type of semantic perturbation, $b$ is the number of perturbations and $\mathcal{V}^{b}_{t}(s)$ is the embeddings of the set of semantic perturbations $\mathcal{A}^{b}_{t}(s)$ around $s$ generated using $\mathcal{P}_{t}$, as described in Section 3.4.
Note that $\mathcal{P}_{t}$ can be given by a collection of different perturbation algorithms and their kinds. The key assumption is that $\mathcal{A}^{b}_{t}(s)$ contains valid sentences semantically similar to $s$ and belonging to the same class. Assuming that membership of $\mathcal{S}$ is easy to compute, then this metric is also easy to compute as the set $\mathcal{A}^{b}_{t}(s)$ is finite and of size $b$, and therefore so is $\mathcal{V}^{b}_{t}(s)$. Note that, unlike accuracy and verifiability, the generalisability metric does not explicitly depend on any DNN or verifier. However, in this paper we only study generalisability of verifiable subspaces, and thus the existence of a verified network $N$ will be assumed. Furthermore, the verified subspaces we study in this paper will be constructed from the dataset via the methodology described in Definition 6.
4.2 Baseline Experiments for Understanding the Properties of Embedding Spaces
The methodology defined thus far has given basic intuitions about the modular nature of the NLP verification pipeline. Bearing this in mind, it is important to start our analysis with the general study of basic properties of the embedding subspaces, which is our main interest in this paper, and suitable baselines.
Benchmark datasets will be abbreviated as “RUAR” and “Medical”. We use $\mathcal{Y}^{pos}$ to refer to the set of sentences in the training dataset with a positive class (i.e. a question asking the identity of the model, and a medical query respectively), and $\mathcal{Y}^{neg}$ to refer to the remaining sentences. For a benchmark network $N:^{m}\rightarrow^{2}$, we train a medium-sized fully-connected DNN (with 2 layers of size (128, 2) and input size 30) using stochastic gradient descent and cross-entropy loss. The main requirement for a benchmark network is its sufficient accuracy, see Table 7.
$N_{base}$ No 93.87 ± 0.14% 93.57 ± 0.18% 96.32 ± 0.05% 94.49 ± 0.26%
Table 7: Mean and standard deviation of the accuracy of the baseline DNN on the RUAR and the Medical datasets. All experiments are replicated five times.
For the choice of benchmark subspaces, we use the following two extreme sets of geometric subspaces:
  1. 1.
    the singleton set containing the maximal subspace $SH(\mathbb{H}(\mathcal{Y}^{pos}))$ around all embedded sentences of the positive class $pos$ in $\mathcal{Y}$. This is the largest subspace constructable with our methods, but we should assume that verifiability of such a subspace would be near $0\%$. It is illustrated in the first graph of Figure 3.
  2. 2.
[…]
We first seek to understand the geometric properties (e.g. volume, $\epsilon$ values) and verifiability figures for these two extremes.
4.3 Verifiability-Generalisability Trade-off for Geometric Subspaces
The number and average volume of the hyper-rectangles that will make up our verified subspaces are shown in Table 8. Generally, we use the following naming convention for our experiments: $\mathbb{H}_{m}$ denotes a hyper-rectangle obtained using a method $m$. For example, RUAR dataset contains $3400$ sentences of the positive class, and therefore the experiment $\mathbb{H}_{\epsilon=0.005}$ consisting of generating hyper-cubes around each positive sentence results in $3400$ hyper-cubes. Using clustering, we obtain a set of $50$, $100$, $200$, $250$ clusters denoted as $\mathbb{H}_{50}$$\mathbb{H}_{250}$ and using the shrinking algorithm we obtain $\mathbb{H}_{sh}$.
Notice the consistent reduction of volume in Table 8, from $\mathbb{H}_{sh}$ to $\mathbb{H}_{50}$ $\mathbb{H}_{250}$ and ultimately to $\mathbb{H}_{\epsilon=0.005}$. There are several orders of magnitude between the largest and the smallest subspace.
Experiment name Hyper-rectangles construction method Avg. volume of hyper-rectangles RUAR Number of hyper-rectangles RUAR Avg. volume of hyper-rectangles Medical Number of hyper-rectangles Medical
$\mathbb{H}_{sh}$ Hyper-rectangle around the entire dataset shrunk to exclude all negative examples - $SH(\mathbb{H}(\mathcal{Y}^{pos}),\mathcal{Y},c^{pos})$ 7.55e-11 1 2.60e-09 1
$\mathbb{H}_{50}$ Set of hyper-rectangles on the dataset separated into 50 clusters - $\mathbb{H}(CL(\mathcal{Y}^{pos},50))$ 1.02e-16 50 6.56e-15 50
$\mathbb{H}_{100}$ Set of hyper-rectangles on the dataset separated into 100 clusters - $\mathbb{H}(CL(\mathcal{Y}^{pos},100))$ 6.23e-18 100 3.25e-17 100
$\mathbb{H}_{200}$ Set of hyper-rectangles on the dataset separated into 200 clusters - $\mathbb{H}(CL(\mathcal{Y}^{pos},200))$ 3.31e-20 200 4.67e-19 200
$\mathbb{H}_{250}$ Set of hyper-rectangles on the dataset separated into 250 clusters - $\mathbb{H}(CL(\mathcal{Y}^{pos},250))$ 6.42e-22 250 2.42e-20 250